### v1

# Qumode Tensor Networks for False Vacuum Decay in Quantum Field Theory

Abstract
False vacuum decay in scalar quantum field theory (QFT) plays a central role in cosmology and particle physics but has remained intractable for classical simulation due to its non-perturbative, highly entangled dynamics. We present a novel method for Hamiltonian simulation of false-vacuum decay in a self-interacting scalar QFT, based on a spatial discretisation into a lattice of bosonic modes. Each mode has a continuous or quasi-continuous local Hilbert space whose dimension is large enough to capture the self-interactions in the Hamiltonian. We refer to these local modes generically as “qumodes”, echoing the ideal case in which the lattice would be realised on a continuous-variable quantum computer. Here we show that the same framework can be implemented on a tensor network, in which case the qumodes are quasi-continuous, corresponding to the local sites of the network having a very high dimensional Hilbert space. Using a time-evolving block decimation algorithm with Trotterised time evolution, we capture the real-time dynamics of the scalar field. In (1+1)-dimensions, we initialise a metastable vacuum and observe the coherent entanglement growth that seeds true-vacuum bubble nucleation. Our approach establishes the qumode network as a scalable framework for non-equilibrium scalar QFT phenomena and paves the way for higher-dimensional studies and continuous-variable quantum computing implementations.
preprint: IPPP/25/37
Figure 1: Schematic overview of simulating quantum field theories with the qumode lattice. A (1+1)-dimensional quantum field theory, with field dynamics governed by a local potential $V(\phi)$, is A]1 discretised in space to form a lattice of qumodes, where each site encodes the field and its conjugate momentum as continuous quadratures $\hat{q}$ and $\hat{p}$, respectively. This qumode lattice can be simulated in two ways: A]A directly on a continuous-variable quantum computer, as outlined in Ref. [1]; or A]2 emulated classically using a qumode TEBD tensor network, where time evolution is approximated by alternating layers of single- and two-mode gates.
I Introduction
Tunnelling in scalar QFT is a cornerstone in our understanding of the dynamics of phase transitions. These processes have profound implications for cosmology, particle physics, and the evolution of the early universe. Indeed, it is thought that they played a central role in setting the initial conditions of the universe, influencing both the large-scale structure of the cosmos and the distribution of matter. Moreover, at least one such transition, namely the electroweak phase transition, is known to have taken place in the early universe. This event could provide explanations for the observed matter-antimatter asymmetry and could generate observable gravitational wave signatures. Thus, these mechanisms provide a crucial theoretical bridge linking microphysical interactions and cosmological scales, and understanding them is essential for reconstructing the history of the cosmos.
At its core, a phase transition in scalar QFT involves the probabilistic escape of the quantum field from a metastable state, the false vacuum, to a lower energy configuration, the true vacuum, by tunnelling through an energy barrier in the potential. This leads to the nucleation of spherical bubbles of true vacuum, which then grow to fill the entire system as the false vacuum “boils off”. The study of this process was pioneered in the seminal works of Callan, Coleman and de Luccia [2, 3, 4], who provided a robust semi-analytical framework for the process. Their analysis focussed on what happens during the nucleation of a single bubble, using the spherical symmetry to reduce the vacuum decay to a one-dimensional quantum mechanical problem, in which the transition pathway between vacua is modelled through the so-called bounce solution. From the perspective of four-dimensional space-time this solution corresponds to the instantaneous production of an extended “lump” of the quantum field known as an instanton, which essentially seeds the nucleating bubble of true vacuum.
This seminal work provided a conceptual framework for analysing the rate of bubble nucleation and it forms the foundation of virtually all the subsequent work that invokes phase transitions in scalar QFT. However, the picture it paints also tells us why in the intervening decades it has never been possible to simulate phase transitions in scalar QFT. The source of the difficulty is the following fact: despite tunnelling being a quantum phenomenon, the large instanton configuration which drives the whole process can somewhat counterintuitively be thought of as a classical object. It is “classical” in the sense that the bounce solution is the exponentially dominant contribution to the tunnelling process, which merely gets dressed by quantum fluctuations. However, this classical lump straddles the false and true vacua and does not belong in either. Consequently, from the perspective of the system sitting in the false vacuum, the instanton configuration that it needs to reach in order to tunnel represents a large coherent state with a huge degree of entanglement. Indeed, thanks to the symmetry of instanton solutions, we can state quite categorically that the longer the lifetime of the false vacuum, the greater the degree of entanglement required to describe its decay. This statement can be formulated more precisely: the entanglement entropy of an instanton configuration is expected to obey an area $\Leftrightarrow$ entropy law, $S_{\text{ent}}\sim\frac{A}{a^{d-2}}$, where $A$ is the surface area of the solution and $a$ is some kind of short-distance cut-off. But the action of the bounce-solution also depends on $A$ as $S\sim\sigma\times A$ where $\sigma$ is the surface tension, such that the rate of bubble formation is proportional to $\Gamma\propto e^{-\sigma A}$.
In summary, then, quantum entanglement is integral to tunnelling in scalar QFT because it underpins the coherence of quantum states across spatial regions. This coherence is essential for the collective behaviour that characterises the formation of nucleating bubbles, and yet it also makes simulating tunnelling in scalar QFT very challenging: one must effectively remain in the false vacuum while generating sufficient entanglement to produce a coherent instanton state. There have been several studies which have sought to simulate aspects of the tunnelling process, which have provided valuable insights [5, 6, 7, 8, 9, 10, 11, 12, 13, 14]. However, in large part due to this difficulty, they considered restricted setups, such as Ising spin-chain models, rather than a full QFT framework.
In this paper, we present an approach that, for the first time, enables the simulation of false vacuum decay within a full scalar QFT framework. Our approach utilises the Hamiltonian formulation of QFT, which bypasses the notorious “sign problem” that is ubiquitous in Monte Carlo techniques for solving the path integral [15, 16, 17]. One key advantage of the Hamiltonian perspective is its natural compatibility with real-time evolution, enabling the study of inherently dynamical processes [18, 19, 20, 8, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 1, 33, 34, 35].
The Hamiltonian formulation has already proven fruitful in studying a wide array of non-equilibrium phenomena, such as quantum quenches [36, 37, 38], thermalisation [39, 40, 41, 42], and dynamic phase transitions in gauge theory [43, 44, 45]. Moreover, as emphasised in foundational work on Hamiltonian quantum simulation [46, 47], these methods offer a scalable route to accessing regimes far beyond current Euclidean techniques [48, 49, 50].
The crucial advance that we will make here is to build a scalar QFT out of a lattice of so-called qumodes, and then leverage the Hamiltonian formalism to simulate its evolution. A qumode is a large-dimensional Hilbert space that describes the quantum mechanics at a local site in space, which crucially must admit arbitrary potentials (Equivalently, it can be thought of as an arbitrary zero-dimensional field theory). By coupling such qumodes together into a lattice, we can build an entire scalar QFT that has an arbitrary potential of our choosing. This allows us to set-up and study the systems discussed above in which the scalar QFT is in a long-lived metastable state that decays to the true vacuum via bubble nucleation.
The procedure is schematically outlined in Fig. 1. We start with the (1+1)-dimensional scalar QFT in the continuum limit, where the dynamics of the quantum field $\phi(x)$ are governed by a Hamiltonian with a free-theory contribution that is quadratic in $\phi(x)$, and a potential, $V(\phi)$. To facilitate the real-time evolution of the field, we A]1 discretise space and place the QFT on a spatial lattice of $N$-sites. The QFT is consequently represented as a system of coupled qumodes, with each qumode associated to a lattice site at position $x_{n}=na$, where $a$ is the lattice spacing. In this qumode-lattice formulation, spatial derivatives are replaced by finite differences, resulting in nearest-neighbour couplings between qumodes, and the emergence of an effective local-potential acting on individual sites [24, 1].
The ideal system to realise such a configuration would be A]A a continuous-variable quantum computing device (CVQC), in particular a photonic platform [51, 52, 53]. CVQC devices offer natural access to bosonic degrees of freedom, long coherence times, and room-temperature operation [54, 55, 56], making them especially well suited for simulating scalar field dynamics. In principle on such devices the dimensionality of the Hilbert space of a single qumode would be infinite. Thus these platforms would provide a promising route toward scalable simulations of vacuum decay and related non-perturbative quantum processes in scalar QFT.
Unfortunately, CVQC devices are not yet large enough to construct a lattice of any great size. Therefore in this paper we demonstrate the method using A]2 a qumode tensor network, by identifying a further possible approximation, which can be achieved by truncating the Hilbert space of each of the qumodes to a finite dimension, $d$, such that qumode operations now become $(d\times d)$ matrices. Therefore, to simulate the real-time dynamics of the QFT, we construct a Time-Evolving Block Decimation (TEBD) algorithm [57, 58]. The TEBD algorithm represents the many-body wavefunction as a Matrix Product State (MPS), which in effect compresses the expoentially large Hilbert-space amplitudes into a sequence of tensors whose bond dimensions reflect the amount of quantum entanglement. The time evolution of the nearest-neightbour Hamiltonian is then performed using the Suzuki-Trotter decomposition [59, 60], where the full time-evolution operator,
$U(t)=e^{-iHt}\leavevmode\nobreak\ ,$ (1)
is approximated with
$\mathcal{U}(t)=\left[\prod_{i}e^{-iH_{i}\delta t}\right]^{t/\delta t}\leavevmode\nobreak\ ,$ (2)
where the Hamiltonian has been split up into a sum of non-commuting parts, $H=\sum_{i}H_{i}$. The total evolution time has been Trotterised by discretising $t$ into small steps of $\delta t$.
The TEBD algorithm implements the Trotterised time-evolution as a series of single-site gate operations, corresponding to the diagonal, on-site contributions from the Hamiltonian, and two-site gate operations, corresponding to the nearest-neighbour interaction contribution to the Hamiltonian. After each Trotter-step of time $\delta t$, the MPS is truncated to a manageable bond dimension. This process is shown for a single Trotter step in routine A]2 of Fig. 1.
II Qumode Lattice QFT Framework
As outlined in the Introduction we shall in this paper be interested in constructing scalar QFTs with potentials that make them unstable to decay. A proto-typical class of such theories are renormalizable scalar field theories with potentials of the following form:
${\mathscr{V}}(\varphi)\leavevmode\nobreak\ =\leavevmode\nobreak\ \frac{\lambda}{4!}(\varphi^{2}-\varphi_{0}^{2})^{2}-\varepsilon\varphi\leavevmode\nobreak\ -\leavevmode\nobreak\ {\mathscr{V}}_{0}\leavevmode\nobreak\ ,$ (3)
where $\lambda$, $\varepsilon$ and ${\mathscr{V}}_{0}$ are constants. The important feature of the potential is that it has a metastable false vacuum at $\varphi_{\rm fv}\approx-\varphi_{0}$. This is where we shall start the field theory. We show a concrete example with $\lambda=0.5$, $\varepsilon=0.1$ and ${\mathscr{V}}_{0}=0.2$, which gives the profile in Fig. 2. The field theory is expected to be able to then quantum tunnel to the true vacuum at $\varphi_{\rm tv}\approx\varphi_{0}$ which is introduced by the second term in the potential. Note that we choose the constant ${\mathscr{V}}_{0}\approx\varphi_{0}\varepsilon$ such that the false minimum at $\varphi_{\rm fv}\approx-\varphi_{0}$ has zero energy, while the true minimum at $\varphi_{\rm tv}\approx\varphi_{0}$ has
${\mathscr{V}}(\varphi_{0})\leavevmode\nobreak\ \approx\leavevmode\nobreak\ -2\varphi_{0}\varepsilon\leavevmode\nobreak\ .$ (4)
Figure 2: The polynomial potential of Eq. (3) with $\varphi_{0}=2$, $\lambda=0.5$, $\varepsilon=0.1$ and ${\mathscr{V}}_{0}=0.2$.
II.1 Qumode preparation
Let us begin by constructing the individual qumodes that will comprise the lattice. As described in App. A we use the exact correspondence between scalar free field theory and a coupled lattice of qumodes in the infinite lattice limit. Thus the entire field theory Hamiltonian with a lattice of $N$ qumodes is given by
$H_{\rm QFT}\leavevmode\nobreak\ =\leavevmode\nobreak\ \frac{a}{2}\sum_{n=1}^{N}\left[{\hat{p}}_{n}(t)^{2}+\left(\frac{\hat{q}_{\,{n+1}}(t)-\hat{q}_{n}(t)}{a}\right)^{2}+2\mathscr{V}(\hat{q}_{n})\right]\leavevmode\nobreak\ ,$ (5)
for arbitrary QFT potential $\mathscr{V}(\varphi)$. Note that we do not impose periodic boundary conditions as these are not straightforward to implement on the GPU systems that we will ultimately be using: we therefore anticipate edge effects which we will have to live with. The local qumode Hamiltonian at space-position $n$ is given by extracting the diagonal parts of the total Hamiltonian:
$H_{n}\leavevmode\nobreak\ =\leavevmode\nobreak\ \frac{a}{2}{\hat{p}}_{n}^{2}+a^{-1}{\hat{q}}_{n}^{2}+a\mathscr{V}(\hat{q}_{n})\leavevmode\nobreak\ .$ (6)
We henceforth take lattice spacing $a=1$.
It is useful to study an individual qumode of the system entirely independently as a quantum mechanical system in its own right before we tie the qumodes together into a lattice. This is a useful check of, for example, the time evolution procedure that we will be using. To do this we will consider the diagonal Hamiltonian in Eq. (6) but with the diagonal pieces from the $\nabla\varphi^{2}$ term subtracted: that is we will consider
$\widetilde{H}_{n}\leavevmode\nobreak\ =\leavevmode\nobreak\ \frac{1}{2}{\hat{p}}_{n}^{2}+\mathscr{V}(\hat{q}_{n})\leavevmode\nobreak\ .$ (7)
Studying an individual qumode with this Hamiltonian will allow us to anticipate the behaviour of the full QFT, and to ensure that we are taking a sufficiently large dimensionality of the local Hilbert space. Indeed, the evolution of this individual qumode should essentially correspond to the behaviour of the entire QFT when it is oscillating in a space-independent coherent manner (as the gradient term is zero). Thus we can, for example, determine the maximum lifetime of the QFT in the false vacuum, which would correspond to the time for this individual qumode to tunnel.
In order to construct a qumode we will discretise its local Hilbert space into a large $d$-dimensional space $q_{n}^{j=1\ldots d}$, where we consider eigenvalues $q_{n}^{j}\in[-L,L]$ with
$q_{n}^{j}\leavevmode\nobreak\ =\leavevmode\nobreak\ -L+j\times\delta q$ (8)
where $\delta q=2L/d$. Thus the quantum mechanical wavefunction is given in the $|q_{n}^{i}\rangle$ basis in the obvious way:
$|\psi_{n}\rangle\leavevmode\nobreak\ =\leavevmode\nobreak\ \sum_{i=1}^{d}|q^{i}_{n}\rangle\left\langle q^{i}_{n}\mid\psi_{n}\right\rangle\leavevmode\nobreak\ .$ (9)
The operator $\hat{q}_{n}$ acts diagonally on the wavefunction in this basis, while the operator $\hat{p}^{2}$ is an operator on the site $n$ given by finite difference in the discrete $|q_{n}^{i}\rangle$ basis. Specifically, we have
$\hat{q}_{n}\leavevmode\nobreak$ $=\leavevmode\nobreak\ \begin{pmatrix}q^{1}_{n}&0&\ldots\\ 0&q^{j}_{n}&\\ \vdots&&q^{d}_{n}\end{pmatrix}\leavevmode\nobreak$
$\leavevmode\nobreak\ \leavevmode\nobreak\ \hat{p}_{n}^{2}\leavevmode\nobreak$ $=\leavevmode\nobreak\ -\frac{\hbar^{2}}{\delta q^{2}}\begin{pmatrix}2&-1&\leavevmode\nobreak\ &\leavevmode\nobreak\ \leavevmode\nobreak\ &\ldots\\ -1&2&-1&\leavevmode\nobreak\ &\ldots\\ \leavevmode\nobreak\ &-1&2&\leavevmode\nobreak\ &\ldots\\ \vdots&\leavevmode\nobreak\ &\leavevmode\nobreak\ &\ddots&-1\\ \vdots&\leavevmode\nobreak\ &\leavevmode\nobreak\ &\leavevmode\nobreak\ \leavevmode\nobreak\ -1&2\end{pmatrix}$ (10)
acting on the wavefunction coefficients $\left\langle q_{n}^{j}\mid\psi_{n}\right\rangle$. We will henceforth choose units such that $\hbar=1$. Of course, if one had access to genuine qumodes with infinite $d$ then these operators would be replaced with the corresponding continuous quadrature variable gates (for example the $\hat{p}_{n}^{2}$ operator would be invoked as part of a “rotation” gate as described in Ref. [24]).
The Hamiltonian evolution for this single qumode may be simulated via the Schrödinger time-evolution operator,
$\widetilde{\mathcal{U}}_{n}(t)\leavevmode\nobreak\ =\leavevmode\nobreak\ e^{-i\widetilde{H}_{n}t}\leavevmode\nobreak\ ,$ (11)
which propagates an initial state $|\psi_{n}(t=0)\rangle$ to the state $|\psi_{n}(t)\rangle=\widetilde{\mathcal{U}}_{n}(t)|\psi(0)\rangle$ at time $t$. However it is convenient to also split the $\hat{p}_{n}$ and $\hat{q}_{n}$ parts, so that the evolution operator is approximated by the following Trotter-Suzuki decomposition [59, 60]:
$\widetilde{\mathcal{U}}_{n}(t)\leavevmode\nobreak\ =\leavevmode\nobreak\ \left[e^{-i\hat{p}_{n}^{2}\delta t}e^{-i\mathscr{V}(\hat{q}_{n})\delta t}\right]^{t/\delta t}+\mathcal{O}\left(\delta t^{2}\right)\leavevmode\nobreak\ .$ (12)
This Trotterisation method approximates the time-evolution operator $\widetilde{\mathcal{U}}_{n}(t)$ up to an error of $\mathcal{O}\left(\delta t^{2}\right)$, and thus is a good approximation for $\delta t\ll 1$. Trotterisation will be unavoidable later when we build the full QFT, thus using it now on a single qumode is a good opportunity to test the quality of the Trotter-Suzuki approximation in the regime of interest.
In order to test the evolution we start the qumode in the ground state of the SHO around the false vacuum at $\varphi=\varphi_{\rm fv}=-2$. The evolution of a single qumode from this initial state is shown in Fig. 3, where in the first panel we show the evolution of the wavefunction of the continuous qumode for times up to $t=60$, and in the second panel we show the evolution in the discretised qumode with a $d=50$ dimensional Hilbert space. The Trotterised time-evolution has been performed with a time step of $\delta t=0.05$. As we will ultimately be interested in the local field expectation value, $\langle\varphi(x)\rangle\equiv\langle\hat{q}_{n}\rangle$, this approximation is more than adequate for our purposes. (Of course it can be improved by taking smaller Trotter steps and higher dimensionality of local Hilbert spaces.) More generally we see that the evolution displays the expected tunnelling behaviour. But what we also observe is that the tunnelling does not “complete” because, in what is effectively a (1+0)-dimensional field theory, the excess energy has nowhere to dissipate. As we shall see, once the qumodes are tied together into a lattice, this excess energy will ultimately be spent driving bubble expansion.
Figure 3: Exact evolution of a single qumode under the potential in Fig. 2 (first panel) versus a Trotterised evolution (second panel) in which the local Hilbert space is discretised as 50-dimensional, while the Trotter time-step is $\delta t=0.05$. Thus, the evolution represents up to $1200$ Trotter steps, respectively, for the five plots.
II.2 Qumode initialisation by adiabatic evolution
Now let us finesse the desired starting state for a qumode that is sitting in the false vacuum. Guided by the usual field theory discussions, one might be inclined as above to use the Simple Harmonic Oscillator (SHO) ground state of the potential expanded around the metastable minimum up to quadratic terms, with cubic and higher terms neglected. That is, if we trivially rewrite the quartic potential as an SHO potential around the false vacuum plus remaining terms as follows,
${\mathscr{V}}(\varphi)\leavevmode\nobreak\ =\leavevmode\nobreak\ \frac{\omega^{2}}{2}(\varphi-\varphi_{\rm fv})^{2}+{\mathscr{V}}_{1}(\varphi)$ (13)
where in the quartic case at hand we have $\varphi_{\rm fv}\approx-\varphi_{0}$, and hence
$\omega^{2}\leavevmode\nobreak\ =\leavevmode\nobreak\ \frac{\lambda\varphi_{0}^{2}}{3}\leavevmode\nobreak\ ,$ (14)
and
${\mathscr{V}}_{1}(\varphi)\leavevmode\nobreak\ =\leavevmode\nobreak\ \frac{\lambda}{4!}(\varphi+\varphi_{0})^{3}(\varphi-3\varphi_{0})-\varepsilon\varphi\leavevmode\nobreak\ -\leavevmode\nobreak\ {\mathscr{V}}_{0}\leavevmode\nobreak\ .$ (15)
Thus, a first approximation to the desired metastable state is the ground state of the SHO with mass unity and frequency $\omega$, namely
$\langle q_{n}|\psi_{n}(s=0)\rangle\leavevmode\nobreak\ =\leavevmode\nobreak\ \left(\frac{\omega}{\pi}\right)^{1/4}e^{-\omega(q_{n}+\varphi_{0})^{2}/2}\leavevmode\nobreak\ .$ (16)
Unfortunately, this approximation is a little too crude for our purposes. Indeed, inspecting the potential of Fig. 2 we see that the potential well around the false vacuum at $\varphi\approx-\varphi_{0}$ is, unlike the SHO potential, noticeably asymmetric. As a result, a field theory that begins with all its qumodes in this SHO ground state will oscillate due to the excess energy. This can be seen in the example shown in Fig. 6. This qumode was initialised in the SHO ground state at $\varphi=-2$ and is “stuck” but continues to oscillate. Moreover, this non-zero energy could significantly change the metastability properties when we come to construct the field theory. Furthermore one may be interested in potentials for which higher order terms dominate over the mass. For most potentials a more precise initial state for the qumodes cannot easily be determined analytically. However it can be determined using adiabatic quantum computing as suggested in  [61, 62, 63, 64, 65].
Let us see how this can work for a single qumode. The aim is to evolve the potential from the SHO potential to an approximation to the full potential according to
${\mathscr{V}}_{s}(\hat{q}_{n})\leavevmode\nobreak\ =\leavevmode\nobreak\ \frac{\omega^{2}}{2}(\hat{q}_{n}-\varphi_{\rm fv})^{2}+s\,({\mathscr{V}}_{1}(\hat{q}_{n})+{\mathscr{V}}_{\rm lift}(\hat{q}_{n}))\leavevmode\nobreak\ ,$ (17)
where $s\in[0,1]$ is an adiabatic evolution parameter, and where ${\mathscr{V}}_{\rm lift}(\hat{q}_{n})$ is a suitable potential that lifts the global minimum at $\hat{q}=\varphi_{\rm tv}$ but does not disrupt the false minimum: for the quartic potential, we can choose
${\mathscr{V}}_{\rm lift}(\hat{q}_{n}))\leavevmode\nobreak\ =\leavevmode\nobreak\ (2+\delta)\,\varphi_{0}\varepsilon e^{-(\hat{q}-\varphi_{0})^{2}}\leavevmode\nobreak\ ,$ (18)
which lifts the minimum at $\hat{q}=\varphi_{0}$ to $+\delta$. Sending the evolution parameter $s$ to one gradually reintroduces all of the remaining ${\mathscr{V}}_{1}$ terms in the potential around the false vacuum at $\hat{q}=\varphi_{\rm fv}=-\varphi_{0}$, while the lift potential ${\mathscr{V}}_{\rm lift}$ prevents the true minimum of the full potential appearing at $\hat{q}_{n}=\varphi_{\rm tv}=+\varphi_{0}$. According to the adiabatic theorem, the wavefunction will remain in the ground state throughout, provided that $s$ evolves sufficiently slowly. The lift potential ensures that the wavefunction does not simply end up in the ground state of the full potential during this process, which of course is centred around $\hat{q}_{n}=+\varphi_{0}$. A comparison of the three potentials is shown in Fig. 4 for the quite extreme case of $\lambda=1$ and $\varepsilon=0.1$.
Figure 4: Initialising the false vacuum state using adiabatic evolution with $\lambda=1$ and $\varepsilon=0.1$: the solid line is the full potential, the dotted line is the SHO approximation around the false vacuum where we begin the qumode in the corresponding SHO groundstate, and the dashed line is the full potential plus lift potential, where here we have taken $\delta=4$. During the adiabatic evolution the potential is gradually changed from the dotted SHO potential to the dashed potential.
Figure 5: The SHO groundstate versus the true ground state for the full potential, and the “false vacuum ground state”. Both of these states were found by adiabatically evolving from the SHO ground state. As might have been anticipated from the shape of the full potential compared to the SHO potential, the “false vacuum ground state” is shifted to the right and broadened compared to the SHO groundstate.
We can implement this idea using a Trotter-Suzuki decomposition of the adiabatic evolution into steps $\delta s=1/M$, by applying the following gates to the initial qumode:
$\widetilde{\mathcal{W}}_{n}\leavevmode\nobreak\ =\leavevmode\nobreak\ \prod_{r=1}^{M}\left[e^{-i\hat{p}_{n}^{2}\delta t}e^{-i\frac{\omega^{2}}{2}(\hat{q}_{n}-\varphi_{\rm fv})^{2}\delta t}\leavevmode\nobreak\ {\widetilde{\mathcal{V}}}^{\,r}_{n}\right]\leavevmode\nobreak\ ,$ (19)
where
$\widetilde{\mathcal{V}}_{n}\leavevmode\nobreak\ =\leavevmode\nobreak\ e^{-i{\mathscr{V}}_{1}(\hat{q}_{n})\delta s\,\delta t}\leavevmode\nobreak\ ,$ (20)
and where the Trotter step $\delta t$, which is chosen to be appropriately small, incorporates normal time evolution. Note that this requires only the single non-linear gate $\widetilde{\mathcal{V}}_{n}$ for a qumode, and in principle one need only determine and record the correct initial qumode state once for any given potential (although in practice it is possible to simply include this procedure as part of the qumode initialisation of each QFT).
In order for the process to remain adiabatic, the total time $t_{\rm tot}=M\delta t$ must be sufficiently long. The criterion proposed in Ref. [64] is that the evolution satisfies
$t_{\rm tot}\leavevmode\nobreak\ \gg\leavevmode\nobreak\ \frac{1}{(\mathscr{E}_{1}-\mathscr{E}_{0})_{\rm min}^{2}}\leavevmode\nobreak\ ,$ (21)
[…]
and adiabatic evolution through tens of Trotter steps suffices to determine the initial state in most cases.
We show an example of this procedure in Fig. 5 for the quartic potential of Fig. 4 (which is a somewhat extreme case). The curves shown are the initial SHO ground state which is centred around $\varphi=-\varphi_{0}$, the true ground state which is of course centred around $\varphi=\varphi_{0}$, found by adiabatically evolving to the full potential (requiring at least $10^{4}$ Trotter steps in order to remain adiabatic in this case), and the “false vacuum ground state” found by adiabatically evolving to the full $+$ lift potential.
Figure 6: As in Fig. 3 but for a ‘stuck’ single qumode with $\lambda=4$.
II.3 Non-renormalisable theories: the double $\tanh(\varphi^{\ell})$ well potential
Figure 7: The double $\tanh\varphi^{\ell}$ potential of Eq. (24) with $\varphi_{\rm tv}=-\varphi_{\rm fv}=2$, $\ell=4$ and $\varepsilon=0.2$. Potentials such as this one with $\ell>2$ give rise to “massless” QFTs.
It is interesting to consider alternatives to the quartic theories of Eq. (3). Potentials with flat barriers are of interest because they can lead to homogeneous decay processes that are not described by the usual Callan-Coleman instanton (see for example [66] for a review and references thereof). Such theories can, for example, undergo tunnelling through the Fubini instanton (in four dimensions), or can tunnel preferentially via thermal or stochastic processes. They can also give rise to long lived oscillations known as “oscillons” ([67, 68] – see Ref. [69] for a recent review of the phenomenon).
A nice example of this is the following potential:
${\mathscr{V}}(\varphi)\leavevmode\nobreak$ $=\leavevmode\nobreak\ \frac{\lambda}{\ell!}\tanh(\mu^{-\ell}(\varphi-\varphi_{\rm fv})^{\ell})$ (24)
$\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ +\leavevmode\nobreak\ \left(\frac{\lambda}{\ell!}+\varepsilon\right)\left(\tanh(\mu^{-\ell}(\varphi-\varphi_{\rm tv})^{\ell})-1\right)\leavevmode\nobreak\ ,$
which we show in Fig. 7 for $\ell=4$, $\mu=1$ and $\lambda=12$. (Note that we add the $\mu$ coefficients for dimensions.)
These potentials would rise to non-renormalisable QFTs in four space-time dimensions, that is, QFTs that have terms of order higher than $\varphi^{4}$ in their potentials. (Ultimately we have four dimensional space-time in mind despite the fact that for the current study we will limit our discussion to (1+1)-dimensional QFTs). Indeed expanding the hyperbolic functions, we see that the potential around the false minimum is approximated by
${\mathscr{V}}(\varphi)\leavevmode\nobreak$ $=\leavevmode\nobreak\ \frac{\lambda}{\ell!}(\varphi-\varphi_{\rm fv})^{\ell}-\frac{\lambda}{3\ell!}(\varphi-\varphi_{\rm fv})^{3\ell}\leavevmode\nobreak\ +\leavevmode\nobreak\ \ldots$ (25)
We thus see another interesting aspect of such potentials which is that they encompass massless theories, that is, theories that have no $\varphi^{2}$ term in the potential. For example, the renormalisable terms in the $\ell=4$ example define a pure $\lambda\varphi^{4}$ theory which is corrected by non-renormalisable terms only at very high order, $\varphi^{12}$. One expects such “flat potential” theories to behave quite differently from generic massive scalar theories.
The correct “false vacuum ground state” can be determined adiabatically as described above, regardless of the fact that the theory is massless. Indeed applying a lift potential ${\mathscr{V}}_{\rm lift}(\hat{q}_{n})$ in such theories is much simpler because one can choose a ${\mathscr{V}}_{\rm lift}$ which simply removes the second term (which corresponds to the global minimum) from ${\mathscr{V}}$. We show an example in Fig. 8 for the $\ell=4$ theory. In this massless case, there is obviously no preferred mass for the initial SHO potential, and one can simply choose its value so that the initial SHO ground state energy is comparable to that of the final ground state energy to minimise the constraint in Eq. (21).
Figure 8: The SHO groundstate versus the “false vacuum ground state” for the double $\tanh\varphi^{4}$ well potential of Eq. (24), with $\lambda=12$.
II.4 Solvable non-renormalizable qumodes: the Pöschl-Teller potential
It is also interesting to consider theories based on the Pöschl-Teller (PT) potential, shown in Fig. 9. These theories are outliers in quantum mechanics because they display unusual phenomena (see [70] for a review), and they are solvable. This means that in classical simulations one can initialise the qumodes analytically, avoiding the need for any kind of adiabatic initialisation process described above for the quartic potentials.
Figure 9: The double Pöschl-Teller potential of Eq. (26) with $\varphi_{\rm tv}=-\varphi_{\rm fv}=2$, $\alpha=1$, $\gamma=2$ and $\varepsilon=0.5$.
Moreover, they are of interest in the present context because the field theories that they will give rise to are also non-renormalisable. Therefore, we will also consider the potential
${\mathscr{V}}(\varphi)\leavevmode\nobreak$ $=\leavevmode\nobreak\ \frac{\alpha^{2}\gamma(\gamma+1)}{2}\tanh^{2}(\alpha(\varphi-\varphi_{\rm fv}))$ (26)
$\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ +\leavevmode\nobreak\ \left(\frac{\alpha^{2}\gamma(\gamma+1)}{2}+\varepsilon\right){\rm sech}^{2}(\alpha(\varphi-\varphi_{\rm tv}))\leavevmode\nobreak\ .$
For analytical purposes, this potential, which is essentially a double PT potential-well, can be very well approximated using the analytic methods that are applicable to the single PT potential because the two separate minima can be exponentially well separated.
Consider starting the qumode sitting in the ground state of the single PT potential,
${\mathscr{V}}_{\rm PT}(\varphi)\leavevmode\nobreak$ $=\leavevmode\nobreak\ \frac{\alpha^{2}\gamma(\gamma+1)}{2}\tanh^{2}(\alpha(\varphi-\varphi_{\rm fv}))\leavevmode\nobreak\ .$ (27)
The energy eigenstates can be solved exactly by the method of ladder-operators, with the normalised ground state wavefunction being given by
$\psi_{n}^{(0)}(q_{n})\leavevmode\nobreak\ =\leavevmode\nobreak\ \langle q_{n}|0\rangle_{-}\leavevmode\nobreak\ =\leavevmode\nobreak\ {\cal N}_{0}\leavevmode\nobreak\ {\rm sech}^{\gamma}(\alpha(q_{n}-\varphi_{\rm fv}))$ (28)
with energy given by
$\mathscr{E}^{-}_{0}\leavevmode\nobreak\ =\leavevmode\nobreak\ \frac{\alpha^{2}}{2}\gamma\leavevmode\nobreak\ ,$ (29)
and with normalisation constant
${\cal N}_{0}\leavevmode\nobreak\ =\leavevmode\nobreak\ {\sqrt{\frac{\alpha}{\sqrt{\pi}}\frac{\,\Gamma\left(\gamma+\frac{1}{2}\right)}{\leavevmode\nobreak\ \Gamma(\gamma)}}}\leavevmode\nobreak\ .$ (30)
The entire tower of states can be written more generally in terms of associated Legendre polynomials of the form
$\psi_{n}^{(\nu)}(q_{n})\leavevmode\nobreak\ =\leavevmode\nobreak\ \langle q_{n}|\nu\rangle\leavevmode\nobreak\ =\leavevmode\nobreak\ {\cal N}^{\gamma-\nu}_{\gamma}P^{\gamma-\nu}_{\gamma}(\tanh(\alpha(q_{n}-\varphi_{\rm fv})))\leavevmode\nobreak\ ,$ (31)
where ${\cal N}^{m}_{\ell}=\sqrt{{m}{\alpha}\frac{(\ell-m)!}{(\ell+m)!}}$, with $\nu=0\ldots\gamma-1$, and energy
$\mathscr{E}^{-}_{\nu}\leavevmode\nobreak\ =\leavevmode\nobreak\ \frac{\alpha^{2}}{2}\left(\gamma(\gamma+1)-(\gamma-\nu)^{2}\right)\leavevmode\nobreak\ .$ (32)
Note that in the limit $\alpha\to 0$ the vacuum energy becomes the simple harmonic oscillator (SHO) vacuum energy of $1/2$ (which is one way to show that $\lim_{\alpha\to 0}\left[{\rm sech}^{1/\alpha^{2}}(\alpha q_{n})\right]\propto e^{-q_{n}^{2}/2}$).
The same method can be applied to wavefunctions localised around the single PT well that approximates the true minimum at $q_{n}\approx\varphi_{\rm tv}$ with its prefactor $\left(\frac{\alpha^{2}\gamma(\gamma+1)}{2}+\varepsilon\right)$. There is an interesting choice of $\varepsilon$ for this potential, namely $\varepsilon=\alpha^{2}(\gamma+1)$. For this choice the prefactor becomes $\frac{\alpha^{2}(\gamma+1)(\gamma+2)}{2}$, and the energies in the PT well around the true minimum become
$\mathscr{E}^{+}_{\nu}\leavevmode\nobreak$ $=\leavevmode\nobreak\ \frac{\alpha^{2}}{2}\left((\gamma+1)(\gamma+2)-(\gamma+1-\nu)^{2}\right)\leavevmode\nobreak\ -\varepsilon\leavevmode\nobreak$
$=\leavevmode\nobreak\ \frac{\alpha^{2}}{2}\left(\gamma(\gamma+1)-(\gamma+1-\nu)^{2}\right)\leavevmode\nobreak\ \leavevmode\nobreak\ .$ (33)
We thus find that $\mathscr{E}^{-}_{0}=\mathscr{E}^{+}_{1}$ for this choice of $\varepsilon$, so that the metastable ground state around the minimum at $q_{n}=\varphi_{\rm fv}$ has the same energy as the first excited state around the true minimum at $q_{n}=\varphi_{\rm tv}$. In such a situation, the ground state of the PT well at $\varphi=\varphi_{\rm fv}$ is actually a superposition of just the two true energy eigenstates of the whole system that have nearly degenerate energies. If we denote the towers of states in each potential well as $|\nu\rangle_{\pm}$, then we can identify these first two excited energy eigenstates (up to irrelevant phases) as
$|{\mathscr{E}}_{1}\rangle\leavevmode\nobreak$ $=\leavevmode\nobreak\ \cos\theta|0\rangle_{-}+\sin\theta|1\rangle_{+}$
$|{\mathscr{E}}^{\prime}_{1}\rangle\leavevmode\nobreak$ $=\leavevmode\nobreak\ -\sin\theta|0\rangle_{-}+\cos\theta|1\rangle_{+}\leavevmode\nobreak\ ,$ (34)
where ${\mathscr{E}}^{\prime}_{1}\gtrsim{\mathscr{E}}_{1}\approx{\mathscr{E}}^{-}_{0}={\mathscr{E}}^{+}_{1}$, for some angle $\theta$. (As the potential is not symmetric we cannot use parity to fix the coefficients.) Conversely if we start the qumode in the ground state $|0\rangle_{-}$ of the metastable minimum, it will evolve as
$|\psi_{n}(t)\rangle\leavevmode\nobreak$ $=\leavevmode\nobreak\ \cos(\Delta\mathscr{E}\,t/2)\,|0\rangle_{-}$ (35)
$\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ +\leavevmode\nobreak\ i\sin(\Delta\mathscr{E}\,t/2)\,\left(\cos 2\theta|0\rangle_{-}+\sin 2\theta|1\rangle_{+}\right)\leavevmode\nobreak$
where $\Delta\mathscr{E}={\mathscr{E}}^{\prime}_{1}-{\mathscr{E}}_{1}$. Now the energy splitting between these two states is known to be generated non-perturbatively by the ($1+0$ dimensional) tunnelling instanton, and is given by
$\Delta\mathscr{E}\leavevmode\nobreak\ \sim\leavevmode\nobreak\ {\mathscr{E}}_{1}e^{-S_{1}}$ (36)
where the instanton action is
[…]
Given Eq. (35), for this particular choice of $\varepsilon$ we expect the state to oscillate back and forth between the initial state and the first excited state around the true minimum in a manner determined by the mixing angle $\theta$ and the instanton action $S_{1}$.
Figure 10: Trotter-Suzuki evolution of a single qumode under the Pöschl-Teller potential in Eq. (26) with $\alpha=1$, $\varphi_{\rm tv}=-\varphi_{\rm fv}=2$, $\gamma=2$ and $\varepsilon=3$. The dashed line represents the first excited state, $\langle q|1\rangle_{+}\sim P^{1}_{2}(\tanh q)$ of the PT potential-well at $q_{n}=2$.
The numerical evolution is shown in Fig. 10, which clearly conforms to this expectation. Thus this figure constitutes a direct measurement of the instanton in quantum mechanical tunnelling using a single qumode. In fact, since the initial $|0\rangle_{-}$ state oscillates to pure $|1\rangle_{+}$, we can conclude that $\theta=\pi/4$, so that, despite the potential being asymmetric, the energy eigenstates consist of the same odd/even linear superpositions $|0\rangle_{-}\pm|1\rangle_{+}$ that would be found for the symmetric potential. In addition we determine the energy splitting $\Delta{\mathscr{E}}$ that the instanton induces in this pair of states to be $\Delta{\mathscr{E}}=\pi/21$.
II.5 The qumode lattice
Now that we understand how to construct and initialise a single qumode, and hence reproduce their local quantum mechanics, let us tie them together into a one-dimensional lattice to reproduce QFT. First of course we must augment the operator $\widetilde{\mathcal{U}}_{n}$ with the contribution to the Hamiltonian coming from the diagonal terms in Eq.(6). That is we have
${\mathcal{U}}_{n}(\delta t)\leavevmode\nobreak\ =\leavevmode\nobreak\ e^{-i\hat{p}_{n}^{2}\delta t}e^{-i\left(\hat{q}_{n}^{2}+\mathscr{V}(\hat{q}_{n})\right)\delta t}\leavevmode\nobreak\ ,$ (38)
with overall
${\mathcal{U}}_{\rm diag}(\delta t)\leavevmode\nobreak\ =\leavevmode\nobreak\ \prod_{n=1}^{N}{\mathcal{U}}_{n}(\delta t)\leavevmode\nobreak\ .$ (39)
Next we come to the non-diagonal contribution to the QFT evolution which arises from the cross-terms in the discretised Hamiltonian of Eq. (5); namely we are required to also consider the Hubbard-like evolution operator
$\mathcal{U}_{\textrm{hop}}(\delta t)\leavevmode\nobreak\ =\leavevmode\nobreak\ \prod_{n=1}^{N}e^{i\,\hat{q}\,_{{n+1}}\,\hat{q}_{n}\,\delta t}\leavevmode\nobreak$ (40)
acting on the lattice. Thus we have in the first order Trotter-Suzuki decomposition a total evolution operator of the form
${\cal U}(\delta t)\leavevmode\nobreak\ =\leavevmode\nobreak\ \mathcal{U}_{\textrm{diag}}(\delta t)\leavevmode\nobreak\ \mathcal{U}_{\textrm{hop}}(\delta t)$ (41)
with the exact evolution operator over time $t$ being expressed as
${\cal U}_{\rm exact}(t)\leavevmode\nobreak$ $=\leavevmode\nobreak\ \left[{\cal U}(\delta t)\right]^{t/\delta t}+{\cal O}(\delta t^{2})$ (42)
It is possible to make improvements (at the expense of adding more gates) which make the Trotter-Suzuki decomposition effectively higher order. In this study we will use instead the second order decomposition
${\cal U}(\delta t)\leavevmode\nobreak\ =\leavevmode\nobreak\ \mathcal{U}_{\textrm{diag}}(\delta t/2)\leavevmode\nobreak\ \mathcal{U}_{\textrm{hop}}(\delta t/2)\leavevmode\nobreak\ \mathcal{U}^{T}_{\textrm{hop}}(\delta t/2)\leavevmode\nobreak\ \mathcal{U}_{\textrm{diag}}(\delta t/2)\leavevmode\nobreak\ ,$ (43)
which gives
${\cal U}_{\rm exact}(t)\leavevmode\nobreak$ $=\leavevmode\nobreak\ \left[{\cal U}(\delta t)\right]^{t/\delta t}+{\cal O}(\delta t^{3})\leavevmode\nobreak\ .$ (44)
Note that there is a trade-off to be made. It is computationally twice as expensive to construct the gates required for such a second order evolution. However, this is a one-time calculation that must be performed at the beginning of the evolution, and meanwhile the number of time steps required to achieve the same precision in the evolution decreases by a power of $2/3$. For very entangled systems it is likely to be favourable to reduce the number of Trotter steps that have to be evaluated.
Let us now turn to the practical implementation of a qumode lattice on a tensor network (TN) [71]. Using the matrix product state (MPS) [72, 73] and matrix product operator (MPO) [74] ansätze, we simulate the second-order Trotterised real-time evolution of (1+1)-dimensional scalar field theory using the time-evolving block decimation (TEBD) algorithm [57, 58], which proceeds in three main stages.
MPS are a class of variational ansätze that efficiently capture the low-entanglement structure of one-dimensional quantum many-body systems. Formally, an MPS expresses the wavefunction of a system with N sites as a chain of rank-3 tensors $\{A^{[n]}\}$, such that the total state reads
$|\psi\rangle=\sum_{\{i_{n}\}}A^{[1]i_{1}}A^{[2]i_{2}}\cdots A^{[N]i_{N}}|i_{1}i_{2}\ldots i_{N}\rangle,$ (45)
where each tensor $A^{[n]i_{n}}$ carries a physical index $i_{n}$ of dimension $d$, and auxiliary bond indices of dimension $\chi$, the so-called bond dimension. The bond dimension controls the amount of bipartite entanglement that can be encoded across a cut in the system and thus determines the expressiveness of the ansatz: low-$\chi$ MPS efficiently approximate ground and low-energy states of gapped local Hamiltonians due to the area law for entanglement entropy . In the context of time evolution, such as with the TEBD algorithm, the MPS structure allows for efficient application of local gates and entanglement truncation via singular value decomposition, thereby rendering the simulation of real-time dynamics tractable even when the full Hilbert space grows exponentially with system size.
Thus, one must construct a MPS representation of the initial state by factorising the exponentially large Hilbert space of the system into a product state of local Gaussian wavepackets. The MPS is then evolved forward in time by applying one and two-site Trotterised gate operations are implemented as matrix operations of the form shown in Equations (39) and (40), respectively. Finally, we perform an entanglement truncation on the system. This is done by truncation the MPS bond-dimension via singular-value decompositions.
Then, the TEBD algorithm is a real-time simulation technique designed for one-dimensional quantum many-body systems with local (typically nearest-neighbour) interactions, and is especially efficient when the quantum state is represented as an MPS. The essential idea of TEBD is to approximate the full time-evolution operator $U(t)=e^{-iHt}$ by Trotter-Suzuki decomposition, splitting H into a sum of terms that act on disjoint pairs of sites, typically of the form
$H=\sum_{\text{even }n}h_{n,n+1}+\sum_{\text{odd }n}h_{n,n+1},$ (46)
so that exponentials of each commuting set can be applied in parallel. A second-order decomposition, for example, proceeds via
$U(t)\approx\left[\prod_{\text{odd }n}\\ e^{-i{\bf h}\delta t/2}\prod_{\text{even }n}e^{-i{\bf h}\delta t}\prod_{\text{odd }n}e^{-i{\bf h}\delta t/2}\right]^{t/\delta t},$
where ${\bf h}\equiv h_{n,n+1}$ and the operators $e^{-ih_{n,n+1}\delta t}$ are two-site unitaries that act locally on pairs of adjacent sites in the MPS. This structure requires the Hamiltonian to be at most nearest-neighbour coupled, or at least decomposable into local gates.
Each gate is applied sequentially by contracting it into the relevant MPS tensors, followed by a singular value decomposition (SVD) to restore the MPS form. The SVD naturally produces a Schmidt decomposition across the bond, and a truncation is applied by discarding singular values below a fixed threshold or keeping only the top $\chi$ singular values. This is what allows TEBD to dynamically control entanglement growth during evolution, ensuring computational feasibility as long as the entanglement entropy remains moderate.
The TEBD algorithm is particularly well-suited for systems with short-range interactions and limited entanglement, such as gapped ground states and mildly entangled real-time dynamics. In the present context, the qumode lattice Hamiltonian consists of a sum of on-site (diagonal) terms and nearest-neighbour interaction terms, rendering it directly compatible with TEBD’s gate-splitting structure. The algorithm proceeds alternately applying single-site gates (on-site potentials and kinetic energy terms) and two-site gates (spatial derivative couplings) to simulate the full dynamics of the field theory.
II.6 Initial state preparation
Finally, we turn to the initial state preparation for the entire qumode lattice. We have seen that it is possible to determine the “false vacuum ground state” of a single uncoupled anharmonic qumode using the adiabatic theorem. However the ground state of the QFT is not simply the product of ground states of the uncoupled anharmonic qumode oscillators. As discussed in Ref. [1] the QFT ground state approaches the product of uncoupled qumode ground states for modes in the large mass and low momentum limit, but generally it will involve multi-qumode entanglement and hence long-range correlations. This distinction could be ignored in Ref. [1] because that work was concerned only with scattering processes, so it could accordingly focus on scattering of low momentum wave-packets. Unfortunately for tunnelling processes the distinction is important, because the long-range fluctuations that seed the process are absent from the product of qumode groundstates. In other words, quantum tunnelling could occur on the lattice spontaneously even if it is initialised as the product of qumode ground states, but it could only be a much rarer essentially zero momentum coherent transition rather than a localised bubble.
In principle one can initialise the entire lattice using the same adiabatic evolution procedure we used for a single qumode in Sec. II.2.
In detail, we begin the system in an MPS which consists of the product of qumodes all in their “false vacuum ground states” determined as described in II.2. Then we act on the MPS with the following set of $M$ MPOs:
$\widetilde{\mathcal{W}}\leavevmode\nobreak\ =\leavevmode\nobreak\ \bigotimes_{r=1}^{M}\prod_{n=1}^{N}\left[e^{-i\hat{p}_{n}^{2}\delta t}e^{-i({\mathscr{V}}(\hat{q}_{n})+{\mathscr{V}}_{\rm lift}(\hat{q}_{n}))\delta t}\leavevmode\nobreak\ {\widetilde{\mathcal{V}}}^{\,r}_{n}\right]\leavevmode\nobreak\ ,$ (47)
where
$\widetilde{\mathcal{V}}_{n<N}\leavevmode\nobreak\ =\leavevmode\nobreak\ e^{-i(\hat{q}_{n}^{2}-\hat{q}_{n}\hat{q}_{n+1})\delta s\,\delta t}\leavevmode\nobreak\ ,$ (48)
and where again we have $\delta s=1/M$ and the small Trotter step $\delta t$ incorporating normal time evolution. As we do not impose periodic boundary conditions, we can take $\widetilde{\mathcal{V}}_{N}=1$.
The principle is, of course, exactly the same as for the determination of the “false vacuum ground state” of a single qumode. That is, beginning in an MPS which is a product of qumode oscillator ground states, and adiabatically turning on the space-derivative couplings, which include the hopping terms, should result in the system reaching the “false vacuum ground state” of the QFT. Note, for this purpose that we have added ${\mathscr{V}}_{\rm lift}$ to the potential to ensure that the system does not tunnel during the adiabatic evolution while it is being prepared. Note also that one could combine everything into a single adiabatic procedure, starting in an MPS which is a product of SHO ground states centred at $\varphi_{\rm fv}$, and adiabatically turning on both ${\mathscr{V}}_{1}$ and the space-derivative terms at the same time.
This kind of procedure would certainly be the initialisation of choice for a qumode lattice built on a genuine continuous-variable quantum computer. However, on a tensor network this procedure would be quite lengthy to implement, and so we will leave it as an exercise for future work.
The alternative method that we will make use of here is to instead approximate the QFT ground state by sampling. That is, one can generate an ensemble (labelled by $s$) of initial position and momentum fluctuations, $\{\Delta q_{n}^{(s)},\Delta p_{n}^{(s)}\}$, whose distributions incorporate those quantum fluctuations and correlations of the QFT ground state which are missing when we simply start the system as an MPS which is a product of quantum anharmonic oscillators. The derivation of these fluctuations is given in App. B. The upshot is that we can mimick the required long-range correlations by setting the qumodes in their “false vacuum ground states” and then augmenting them by random sampling the classical distributions of position and momentum fluctuations presented in Eq. (96), which we reproduce here for convenience:
$\Delta{q}_{n}\leavevmode\nobreak$ $=\leavevmode\nobreak\ \sum_{\alpha=0}^{N-1}\frac{1}{\sqrt{2N}}\sqrt{\left(\frac{1}{\omega}-\frac{1}{\omega_{\alpha}}\right)}\leavevmode\nobreak\ \cos(2\pi n\alpha/N)\leavevmode\nobreak\ {\eta}_{n,\alpha}\leavevmode\nobreak\ ,$
$\Delta p_{n}\leavevmode\nobreak$ $=\leavevmode\nobreak\ \sum_{\alpha=0}^{N-1}\sqrt{\frac{\omega_{\alpha}-\omega}{2N}}\,\cos(2\pi n\alpha/N)\leavevmode\nobreak\ {\eta}^{\prime}_{n,\alpha}\leavevmode\nobreak\ ,$ (49)
where $\eta_{n}$ and $\eta_{n}^{\prime}$ are standard normal vectors, and the angular frequencies $\omega$ and $\omega_{\alpha}$ are given in Eqs. (3) and (71) respectively. (The shift $\Delta p$ is of course implemented by acting on the wave function of each qumode with the operator $e^{i\Delta p_{n}\,\hat{q}_{n}}$.)
III False vacuum decay in (1+1)-dimensional QFT
III.1 The objective: tunnelling solutions
We now turn to the tunnelling problem that we will focus on for this study. For completeness let us first recap the analytic tunnelling results of Refs. [2, 3, 4]. In order to find the instanton tunnelling solution analytically we are required to solve the system in the Euclidean potential, that is with imaginary time $\tau=it$. The solution is as we have mentioned radially symmetric in Euclidean space-time, that is it has $O(2)$ symmetry, and is hence a function of $r=\sqrt{\tau^{2}+x^{2}}$. Thus, in (1+1)-dimensions the Euclidean equations of motion become
$\partial^{2}_{r}\varphi+r^{-1}\partial_{r}\varphi\leavevmode\nobreak\ =\leavevmode\nobreak\ \partial_{\phi}{\mathscr{V}}\leavevmode\nobreak\ .$ (50)
This now resembles the classical equation of motion of a ball rolling in the inverted potential $-\mathscr{V}$, with $r$ playing the role of time and with the $r^{-1}\partial_{r}\varphi$ term providing a friction that decays with time. Thus, the solution with the correct boundary conditions can be found by starting the ball close to the true vacuum at $\varphi\approx\varphi_{0}$ at ‘time’ $r=0$ (corresponding to the centre of the instanton) and letting it come to a halt at $r\to\infty$ precisely at the false vacuum at $\varphi=-\varphi_{0}$. This requires a fine-tuning of the starting value of $\varphi$, numerically solving the one dimensional differential equation in Eq. (50), and employing the overshoot-undershoot method described in Ref. [2] to satisfy the boundary conditions. The result for the quartic potential in Fig. 2 with $\lambda=0.5$, $\varphi_{0}=2$ and $\varepsilon=0.02$ is the bubble profile shown in Fig. 11. Although it is not possible to find a closed form for the radially symmetric solution, it can be very accurately modelled by the following function:
$\varphi(r)\leavevmode\nobreak\ \approx\leavevmode\nobreak\ -\varphi_{0}\,\tanh\beta(r-r_{0})$ (51)
where in this case $r_{0}=21.8$ and the inverse width of the bubble wall is $\beta=0.4$. The behaviour after the bubble formation is found by going back to Minkowski space time:
$\varphi(x,t)\leavevmode\nobreak\ \approx\leavevmode\nobreak\ -\,\varphi_{0}\,\tanh\beta(\sqrt{x^{2}-t^{2}}-r_{0})\leavevmode\nobreak\ .$ (52)
Thus we see that once formed the bubble expands, with the bubble walls following the hyperbolæ $x^{2}-t^{2}=r^{2}_{0}$ and asymptoting to the lightcones at $x=\pm t$. Outside the wall the field drops exponentially fast to the false vacuum, $\varphi=-\varphi_{0}$. Inside the bubble wall the field is in the true vacuum state with tiny oscillations: indeed here we can approximate
$\varphi(x,t>|x|)\leavevmode\nobreak\ \approx-\varphi_{0}\,(1+2e^{-2\beta r_{0}}\cos{2\beta|t^{2}-x^{2}|})\leavevmode\nobreak\ .$ (53)
What is the typical bubble size, $r_{0}$? To find it when the bubble wall is thin compared to $r_{0}$ (the so-called thin-wall approximation) we can compare the surface tension of the bubble to the vacuum energy contained within its volume. The former can be estimated by integrating the energy across the wall,
[…]
where $S_{1}$ is precisely the Euclidean action of Eq. (37). A critical bubble has a radius $r_{0}\approx r_{c}$ for which the vacuum energy inside the bubble is just enough to off-set the energy of the domain walls. Using Eq. (4) this implies $S_{1}=2\varphi_{0}\varepsilon r_{c}$ hence
$r_{c}\leavevmode\nobreak\ =\leavevmode\nobreak\ \frac{S_{1}}{2\varphi_{0}\varepsilon}\leavevmode\nobreak\ .$ (55)
Once formed with $r_{0}\approx r_{c}$, a bubble expands along the aforementioned hyperbolæ and if it is not impeded by friction it approaches the speed of light, driven by the energy gained from the increasing internal volume of true vacuum.
It is straightforward to see that, in fact, any supercritical bubble that is stationary at $t=0$ follows a hyperbola which is determined by its initial radius. Indeed, if we assume that the bubble profile is symmetric $O(2)$, then it is of the form $\varphi(x,t)=f(r-r_{0})$, where $r_{0}$ is simply the bubble radius at $t=0$.
As mentioned in the introduction the rate of bubble nucleation per unit space-time volume is exponentially sensitive to the Euclidean action associated with the tunnelling solution. Specifically, in the thin-wall approximation it follows the Arrhenius-like behaviour $\Gamma\leavevmode\nobreak\ =\leavevmode\nobreak\ \kappa\,e^{-S_{1}}$, where $\kappa$ is a prefactor that accounts for quantum fluctuations around the instanton solution (which is notoriously difficult to evaluate). This exponential suppression means that even small increases in $S_{1}$ can lead to a dramatic decrease in the nucleation rate. Thus the value of $S_{1}$ effectively encodes how “difficult” it is for the system to tunnel from the false vacuum to the true vacuum, with lower $S_{1}$ corresponding to exponentially higher nucleation probabilities and faster phase transitions.
This then is the paradigm for false vacuum decay that was laid down in Refs. [2, 3]. Before considering the equivalent evolution on our qumode lattice, we should emphasize that there are many subtle issues surrounding the dynamics of these processes which have been discussed subsequently, and which are not fully understood: for example, whether or not the bubble “runs away” to light speed or instead reaches a terminal velocity depends on the friction generated by its interactions with the medium it is travelling through, and is a complicated and model-dependent question (see [75, 76, 77, 78] and for a recent review see [79]). This of course is one of the motivations for studying the process numerically.
Figure 11: Bubble profile for $\lambda=0.5$ and $\varepsilon=0.02$ for which Eq. (55) gives $r_{c}=23$.
III.2 Results: tunnelling in (1+1)-dimensional scalar theory
To validate the model, we simulate the real-time dynamics of the system in two distinct scenarios. First, we consider “pre-formed” bubbles, initialising the system with a true-vacuum bubble with an initial radius $r_{0}$. Second, we study phase transitions in the full system by examining bubble nucleation arising from quantum fluctuations.
III.2.1 Isolated bubbles: expansion and collapse
(a) Subcritical bubble
(b) Supercritical bubble
Figure 12: In 13(a) we show a collapsing subcritical bubble in the Pöschl-Teller potential of Eq. (26) with $\varphi_{\rm tv}=-\varphi_{\rm fv}=2$, $\alpha=1$, $\gamma=3/2$ and $\varepsilon=0.25$. Fig. 13(b) shows an expanding bubble in the potential of Eq. (24) with $\varphi_{\rm tv}=-\varphi_{\rm fv}=2$, $\lambda=1$, and $\varepsilon=0.1$, which shows the bubble walls following a hyperbolic trajectory that reaches a terminal speed of 0.72. Note that we do not impose periodic boundary conditions, and the bubble walls “bounce” off the sides of the lattice elastically.
To begin, we examine the real-time dynamics of pre-formed bubbles with an initial radius, $r_{0}$. Figure 12 presents the results from the qumode tensor network for bubbles of different initial radii. The first scenario, shown in Fig. 13(a), considers the collapse of a subcritical bubble in the Pöschl-Teller potential from Eq. (26) with $\varphi_{\rm tv}=-\varphi_{\rm fv}=2$, $\alpha=1$, $\gamma=3/2$ and $\varepsilon=0.25$. In this example, we observe that the bubble fails to reach the critical radius. There are two interesting aspects of the collapse. The first phenomenon that can be seen has to do with the way the collapse occurs. We can clearly see the emission of several wavepackets from the walls. Averaged over time these emissions would generate friction on the walls. Here, however, we see that such emission (together with momentum conservation) is a necessary component of bubble collapse. This is because in one space-dimension there is no surface tension acting to collapse the bubble. In other words in one space-dimension the energy contained within the bubble walls is independent of the size of the bubble. The second interesting aspect of Fig. 13(a) is related. It is the fact that the final moments of collapse, in which the whole bubble “flops” back to the false vacuum, would correspond to super-luminal wall velocity. This in turn implies that the entanglement on the qumode lattice must spread over distances of at least the radius of the collapsing bubble at that instant in time, in order for the collapse to happen.
Figure 13(b) shows the expansion of a bubble in the potential from Eq. (24) with $\varphi_{\rm tv}=-\varphi_{\rm fv}=2$, $\lambda=1$ and $\varepsilon=0.1$. In contrast with Fig. 13(a), this bubble is supercritical and we observe the phenomena outlined in Sec. III.1. As the bubble’s radius is greater than the critical bubble radius, the walls expand. The bubble expansion is entirely driven by internal pressure caused by the energy gain from increasing the volume of the true vacuum. This acceleration is relatively local, and long-range entanglement plays very little roll in the behaviour. As a result, the walls accelerate along hyperbolic trajectories which asymptote to a constant terminal velocity. However the terminal velocity is subluminal ($v_{\rm term}\approx 0.72$), due to the friction mentioned above.
Note that we do not impose periodic boundary conditions and this results in the bubble wall colliding elastically with the sides of the lattice. We can in principle use this phenomenon to measure the mass of the bubble walls. We observe that after the bounce the walls are projected a distance of $d_{\rm bounce}\approx 20$ into the bulk of the lattice before being repelled back by the pressure, $\varepsilon$. (This relatively short distance also tells us that friction has indeed resulted in a terminal subluminal speed.) We may make an rough estimate for the mass of the wall by neglecting the effect of friction after the bounce and taking a non-relativistic approximation. Equating the initial kinetic energy with the maximum potential energy due to the pressure gives
[…]
III.2.2 Phase transitions
(a) $\varphi_{\rm fv}=-2$, $\varphi_{\rm tv}=0.7$, $\varepsilon=0.5$
(b) $\varphi_{\rm fv}=-2$, $\varphi_{\rm tv}=0.7$, $\varepsilon=0.4$
Figure 13: Phase transitions on the qumode lattice for two different parameter settings.
Now let us turn to phase transitions in the system as a whole. In Fig. 13 we show two examples of phase transitions that were produced with the lattice initialised in the random sampling approximation to the QFT false vacuum that was discussed in Sec. II.6. The potential used for these examples is the $\tanh q^{\ell}$ function of Eq. (24) with $\ell=2$. The true minimum for these examples was chosen to be at $\varphi_{\rm tv}=0.5$ in order to allow for relatively rapid transitions. The figures show the production of bubbles which coalesce to complete the phase transition. The first example with $\varepsilon=0.5$ is very rapid because the instanton is very small (with $r_{c}\approx 2$). The second example has $\varepsilon=0.4$ and the instanton size has grown enough ($r_{c}\approx 5$) to noticeably slow the bubble production, as would be anticipated from the analysis in Sec. III.1. One observation that supports the bubble formation being the result of genuine instanton transitions is that we find that the maximum bond dimension has to exceed a certain threshold in order for this behaviour to appear (if the dimension is set to be too small then there is little tunnelling). Moreover, this threshold appears to grow with the instanton size. In future work, an in-depth study of these relations will be carried out, with a particular focus on confirming the exponential dependence of the decay rate $\Gamma$ on the action.
IV Summary and Conclusions
In this work, we have presented a novel framework for simulating false vacuum decay in scalar QFT using a qumode-based Hamiltonian formulation. Our method has the potential to overcome longstanding challenges associated with simulating highly entangled, non-equilibrium tunnelling dynamics in QFT by leveraging a qumode lattice representation and implementing real-time evolution via tensor network techniques, specifically the time-evolving block decimation algorithm.
Unlike previous approaches, which typically restricted attention to spin chains or Euclidean simulations, our framework, based on Hamiltonian simulation, remains firmly within the QFT setting and directly simulates real-time tunnelling phenomena. The core innovation lies in the use of qumodes-bosonic modes with (quasi-)continuous Hilbert spaces, which enable the construction of scalar field theories with arbitrary potentials. This setup can, in principle, be implemented on continuous-variable quantum computing devices in its presented form. However, to enable simulation on current classical hardware, we truncate the local Hilbert space and employ a matrix product state ansatz for the quantum state, thus achieving scalable simulations with tunable entanglement control.
We demonstrated the validity of our method through a sequence of increasingly complex tests. First, we validated the real-time dynamics of isolated qumodes in double-well potentials, including adiabatic preparation of metastable states and analytic tunnelling in exactly solvable potentials such as the Pöschl-Teller well. We then extended the method to full (1+1)-dimensional QFTs on qumode lattices, capturing bubble nucleation and expansion, and reproducing analytic expectations such as hyperbolic wall trajectories and critical bubble dynamics.
Importantly, we showed that quantum tunnelling emerges in this framework only once sufficient entanglement is captured by the MPS. The bond dimension required increases with instanton size, suggesting a direct numerical signature of the entanglement entropy scaling expected for false vacuum decay. This provides a powerful diagnostic and highlights the necessity of entanglement-aware methods in simulating quantum phase transitions.
Compared to previous literature, our work represents the first demonstration of scalar QFT tunnelling using a real-time evolved Hamiltonian lattice system with quasi-continuous local degrees of freedom. While continuous-variable tensor network methods, such as continuous MPS or cPEPS [80, 81], have been developed for ground-state field theory, and continuous-variable quantum computing proposals exist for QFT [24, 82], no prior work has combined qumode-lattice discretisation with real-time tensor-network evolution to simulate false vacuum decay in QFT. Our method is thus, to the best of our knowledge, the first demonstration of qumode tensor networks applied to non-equilibrium tunnelling dynamics in scalar field theory.
It can open the door to systematic studies of vacuum decay in arbitrary scalar potentials, including non-renormalisable theories and those beyond the reach of semiclassical methods. Moreover, our framework is designed to be portable to CVQC platforms, where, if technically realised at scale and high fidelity, genuine continuous-variable computing would be able to improve on tensor network simulations in scenarios with large entanglement.
Future extensions include implementing entangled initial state preparation via full-lattice adiabatic evolution, exploring multi-field tunnelling landscapes, and applying this method to higher-dimensional QFTs. Our results show the viability of qumode tensor networks as a bridge between classical simulation and future quantum computing applications in non-perturbative quantum field theory.
Acknowledgements
This work was supported by the STFC under the IPPP grant ST/T001011/1. Part of this work used the DiRAC@Durham facility managed by the ICC on behalf of the STFC DiRAC HPC Facility (www.dirac.ac.uk). The equipment was funded by BEIS capital funding via STFC capital grants ST/K00042X/1, ST/P002293/1, ST/R002371/1 and ST/S002502/1, Durham University and STFC operations grant ST/R000832/1.
Appendix A Scalar field theory on a qumode lattice
In this Appendix, we recap the relation between scalar quantum field theories and a lattice system of coupled qumode. For this discussion a “qumode” can be thought of as any quantum mechanical oscillator. For the free-field theory, the qumodes in question will correspond to simple harmonic oscillators (SHOs), while in non-linear systems the qumodes will be quantum mechanical oscillators with more general potentials.
A.1 Free fields on the qumode lattice
Let us begin with the Lagrangian for the system we wish to encode, namely a scalar quantum field, $\varphi(x,t)$. To establish our dictionary, it is convenient to begin with a free field, where $\varphi$ has mass $\omega$ given in natural units ($\hbar=c=1$) such that
$\mathscr{L}\leavevmode\nobreak\ =\leavevmode\nobreak\ \frac{1}{2}\left(\partial^{\mu}\varphi(x,t)\partial_{\mu}\varphi(x,t)-\omega^{2}\varphi(x,t)^{2}\right)\leavevmode\nobreak\ ,$ (57)
with $\varphi$ and $\pi$ obeying the usual equal-time commutation relations. The conjugate momentum associated with the field is
$\pi(x,t)\leavevmode\nobreak\ =\leavevmode\nobreak\ \frac{\partial\mathscr{L}}{\partial(\partial_{0}\varphi(x,t))}\leavevmode\nobreak\ =\leavevmode\nobreak\ \partial_{0}\varphi(x,t)\leavevmode\nobreak\ ,$ (58)
such that the Hamiltonian density is
$\mathscr{H}(x,t)\leavevmode\nobreak\ =\leavevmode\nobreak\ \frac{1}{2}\left(\pi(x,t)^{2}+(\nabla\varphi(x,t))^{2}+\omega^{2}\varphi(x,t)^{2}\right)\leavevmode\nobreak\ .$ (59)
We wish to simulate this system by coupling together a lattice consisting of a large number, $N$, of quantum mechanical oscillators. For simplicity, we shall henceforth focus on a one-dimensional field theory and locate the oscillators at positions $x_{n}=an$ where $n\in[0,N-1]$, and where $a$ is the lattice spacing. We will also enforce periodic boundary conditions throughout, such that $x_{N}=x_{0}$, although this boundary condition will not be consequential in our discussion.
Each quantum-mechanical oscillator has its own quadrature operators that we denote $\hat{q}_{n}$ and $\hat{p}_{n}$ (and can initially at least be assigned its own wavefunction $\psi_{n}(q_{n})\equiv\langle q_{n}|\psi_{n}\rangle$). It is straightforward to verify that there is a configuration of nearest-neighbour couplings of the oscillators that reproduces the free-field theory Hamiltonian, $\mathscr{H}$, in the large $N$ limit. Moreover, the field operator $\varphi(x)$ and its corresponding conjugate momentum operator $\pi(x)$ can be directly equated with the quadrature variables of the oscillators (written as a density) at that position on the lattice, such that one can identify
$\hat{q}_{n}(t)\leavevmode\nobreak\ \leftrightarrow\leavevmode\nobreak\ \varphi(x_{n},t)\leavevmode\nobreak\ ,\qquad\hat{p}_{n}(t)\leavevmode\nobreak\ \leftrightarrow\leavevmode\nobreak\ \pi(x_{n},t)\leavevmode\nobreak\ .$ (60)
To make this identification, we first express the spatial derivatives that appear in the Hamiltonian density of Eq. (59) by finite difference,
$\partial_{i}\varphi(x_{n},t)\leavevmode\nobreak\ \leftrightarrow\leavevmode\nobreak\ \frac{{\hat{q}}_{n+1}(t)-{\hat{q}}_{n}(t)}{a}\leavevmode\nobreak\ .$ (61)
We then define the total Hamiltonian by summing the contributions from the coupled individual qumodes over the entire lattice as follows:
$H\leavevmode\nobreak\ =\leavevmode\nobreak\ \frac{a}{2}\sum_{n=0}^{N-1}\left[{\hat{p}}_{n}(t)^{2}+\left(\frac{\hat{q}_{\,\overline{n+1}}(t)-\hat{q}_{n}(t)}{a}\right)^{2}+\omega^{2}\hat{q}_{n}(t)^{2}\right]\leavevmode\nobreak\ ,$ (62)
with variables normalized such that
$[\hat{q}_{n},\,\hat{p}_{m}]\leavevmode\nobreak\ =\leavevmode\nobreak\ i\delta_{nm}\leavevmode\nobreak\ ,$ (63)
and where periodic boundary conditions can be enforced if desired using
$\overline{n+1}\leavevmode\nobreak\ =\leavevmode\nobreak\ n+1\mod(N)\leavevmode\nobreak\ .$ (64)
The $a$ prefactor in the Hamiltonian ensures that it scales extensively with the size of the lattice, ${\rm Vol}=L=aN$. We can also define the combined wavefunction and the combined set of quadrature variables,
$|\Psi\rangle\leavevmode\nobreak\ =\leavevmode\nobreak\ \otimes_{n}|\psi_{n}\rangle\leavevmode\nobreak\ \leavevmode\nobreak\ ;\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ |{\mathbf{q}}\rangle\leavevmode\nobreak\ =\leavevmode\nobreak\ \otimes_{n}|q_{n}\rangle\leavevmode\nobreak\ .$ (65)
In the ground state of the lattice, all the wave functions will be degenerate, and there will be no contribution to the Hamiltonian from the gradient terms:
[…]
Evidently, $\omega$ is the natural frequency of the lattice of oscillators when the qumodes are all degenerate. The ground state energy of the system is then simply the sum of the ground-state energies of all the qumodes, $H_{0}=\frac{\omega}{2}\,aN$, which is equivalently the volume multiplying the ground-state energy of a single qumode. Perturbing this vacuum in a degenerate manner would result in a qumode lattice wavefunction that continued to factorise into a product of independent qumodes,
$\langle{\mathbf{q}}|\Psi_{\rm degen}\rangle\leavevmode\nobreak\ =\leavevmode\nobreak\ \prod_{n}\langle q_{n}|\psi_{n}\rangle\leavevmode\nobreak\ .$ (67)
Of course, propagating modes then correspond to perturbations of the lattice around such a degenerate ground state. However it is important to realise that a lattice of qumodes in their ground state does not correspond to the traditional QFT vacuum, precisely due to the spatial kinetic terms, but it effectively coincides with the QFT vacuum for non-relativistic modes. We will return to the question of the QFT vacuum in detail later.
To make explicit the relation anticipated in Eq. (60), we must first diagonalise $H$. This is made non-trivial by the hopping terms which provide a “circulant” contribution. Defining $\nu=e^{2\pi i/N}$, the diagonalisation matrix is found to be
$U_{\alpha m}\leavevmode\nobreak\ =\leavevmode\nobreak\ \frac{1}{\sqrt{N}}\nu^{\alpha m}\leavevmode\nobreak\ ,$ (68)
where we use Greek letters $\alpha,\beta,\ldots,\kappa$ to refer to the diagonal basis. In this diagonal basis, the Hamiltonian of Eq. (62) becomes,
$H\leavevmode\nobreak\ =\leavevmode\nobreak\ \frac{a}{2}\sum^{N-1}_{\alpha=0}\left(|\hat{p}_{\alpha}|^{2}+\omega_{\alpha}^{2}|\hat{q}_{\alpha}|^{2}\right)\leavevmode\nobreak\ ,\$ (69)
where
$\hat{p}_{\alpha}\leavevmode\nobreak\ =\leavevmode\nobreak\ \sum_{n}U_{\alpha n}\hat{p}_{n}\leavevmode\nobreak\ \leavevmode\nobreak\ ;\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \hat{q}_{\alpha}\leavevmode\nobreak\ =\leavevmode\nobreak\ \sum_{n}U_{\alpha n}\hat{q}_{n}\leavevmode\nobreak\ ,$ (70)
and where the eigenvalues are
$\omega_{\alpha}\leavevmode\nobreak\ =\leavevmode\nobreak\ \sqrt{\left(\omega^{2}+\frac{4}{a^{2}}\sin^{2}\left(\frac{\pi\alpha}{N}\right)\right)}\leavevmode\nobreak\ .$ (71)
Thus, in the $N\to\infty$ limit we recover the usual interpretation of free-field QFT as a tower of SHO oscillators, which have momentum
[…]
and which obey the relativistic relation $\omega_{\alpha}^{2}=\omega^{2}+k_{\alpha}^{2}$.
Two remarks are in order. First note that in the diagonal basis, $\hat{q}_{\alpha}$ and $\hat{p}_{\alpha}$ are complex, with a phase given by the momentum $k_{\alpha}$, despite the fact that $\hat{q}_{n}$ and $\hat{p}_{n}$ are real. Consequently, each real mode in space is now represented by a pair of complex conjugate modes in momentum (Fourier) space, corresponding to left and right-moving momentum eigenstates. As their sum must be real, we can conclude that $\hat{q}_{\alpha}$ can only be considered independently observable in conjunction with momentum conservation in the QFT.
Our second remark is that it is clear from Eq. (71) that the special relativistic relation $\omega_{\alpha}=\sqrt{\omega^{2}+k_{\alpha}^{2}}$ is recovered only under the assumption that $\pi\alpha/N=ak_{\alpha}/2\ll 1$. We can thus anticipate the possibility of non-relativistic artefacts on a finite lattice when the momenta approach unity in units of $a^{-1}$. As the lattice momentum is really $\tilde{k}_{\alpha}=(2/a)\sin(\pi\alpha/N)$ these artifacts would resemble superluminal states that satisfy $\sin^{-1}(a\tilde{k}_{\alpha}/2)\approx ak_{\alpha}/2$, i.e. they would have $\alpha^{\prime}=\alpha+2\beta N$ where $\beta\in{\mathbb{Z}}$. Likewise, the lattice’s highest physical momentum can be described as $k_{\alpha}\sim 1/a$. Hence the lattice-spacing $a$ defines a natural UV cut-off, and relativistic behaviour can only be modelled for modes that have $\omega a\ll 1$.
We are now able to make the direct identification promised in Eq. (60) between $\varphi(x_{n})$ and the quadrature variable at that lattice site $q_{n}$. To do this, recall that the ladder operators of the qumodes on the lattice are
$\hat{a}_{n}$ $\leavevmode\nobreak\ =\leavevmode\nobreak\ \sqrt{\frac{\omega}{2}}\left(\hat{q}_{n}+\frac{i}{\omega}\hat{p}_{n}\right)\leavevmode\nobreak\ ,$
$\hat{a}^{\dagger}_{n}$ $\leavevmode\nobreak\ =\leavevmode\nobreak\ \sqrt{\frac{\omega}{2}}\left(\hat{q}_{n}-\frac{i}{\omega}\hat{p}_{n}\right)\leavevmode\nobreak\ ,$ (73)
normalized such that $H=\sum_{n}a\omega(\hat{a}^{\dagger}_{n}\hat{a}_{n}+\frac{1}{2})$ when all the qumodes are degenerate. In the diagonal basis these become
$\hat{a}_{\alpha}$ $\leavevmode\nobreak\ =\leavevmode\nobreak\ \sqrt{\frac{\omega_{\alpha}}{2}}\leavevmode\nobreak\ \sum_{n}\left(U_{\alpha n}\,\hat{q}_{n}+\frac{i}{\omega_{\alpha}}U_{\alpha n}\,\hat{p}_{n}\right)\leavevmode\nobreak\ ,$
$\hat{a}_{\alpha}^{\dagger}$ $\leavevmode\nobreak\ =\leavevmode\nobreak\ \sqrt{\frac{\omega_{\alpha}}{2}}\leavevmode\nobreak\ \sum_{n}\left(\hat{q}_{n}\left(U^{\dagger}\right)_{n\alpha}-\frac{i}{\omega_{\alpha}}\hat{p}_{n}\left(U^{\dagger}\right)_{n\alpha}\right)\leavevmode\nobreak\ .$ (74)
The fact that the field operator at position $x_{n}$ is simply the quadrature variable at that point on the lattice can be seen as follows: our claim is that
$\varphi(x_{n})$ $\leavevmode\nobreak\ =\leavevmode\nobreak\ \hat{q}_{n}\leavevmode\nobreak\ =\leavevmode\nobreak\ \frac{1}{\sqrt{2\omega}}(\hat{a}_{n}+\hat{a}_{n}^{\dagger})$
$\leavevmode\nobreak\ =\leavevmode\nobreak\ \sum_{\alpha}\frac{1}{\sqrt{2\omega_{\alpha}}}\left(U_{n\alpha}^{\dagger}\hat{a}_{\alpha}+\hat{a}^{\dagger}_{\alpha}U_{\alpha n}\right)\leavevmode\nobreak\ ,$
$\leavevmode\nobreak\ =\leavevmode\nobreak\ \sum_{\alpha}\frac{1}{\sqrt{2\omega_{\alpha}\,N}}\left(\hat{a}_{\alpha}e^{-i\alpha n}+\hat{a}^{\dagger}_{l}e^{i\alpha n}\right)\leavevmode\nobreak\ ,$
$\leavevmode\nobreak\ =\leavevmode\nobreak\ \sum_{\alpha}\frac{1}{\sqrt{2\omega_{\alpha}\,N}}\left(\hat{a}_{\alpha}e^{-ik_{\alpha}x_{n}}+\hat{a}^{\dagger}_{\alpha}e^{ik_{\alpha}x_{n}}\right)\leavevmode\nobreak\ ,$ (75)
where we use the $N\to\infty$ definition of the momentum in Eq. (72). Similarly, its conjugate momentum is
$\pi(x_{n})$ $\leavevmode\nobreak\ =\leavevmode\nobreak\ \hat{p}_{n}\leavevmode\nobreak\ =\leavevmode\nobreak\ \sum_{\alpha}(-i)\sqrt{\frac{\omega_{\alpha}}{2aN}}\left(\hat{a}_{\alpha}e^{-ik_{\alpha}x_{n}}-\hat{a}^{\dagger}_{\alpha}e^{ik_{\alpha}x_{n}}\right)\leavevmode\nobreak\ .$ (76)
[…]
$\varphi(x)$ $\leavevmode\nobreak\ =\leavevmode\nobreak\ \int\frac{\textrm{d}k}{(2\pi)}\frac{1}{\sqrt{2\omega_{k}}}\left(\hat{a}_{k}e^{ikx}+\hat{a}^{\dagger}_{k}e^{-ikx}\right)\leavevmode\nobreak\ ,$
$\pi(x)$ $\leavevmode\nobreak\ =\leavevmode\nobreak\ \int\frac{\textrm{d}k}{(2\pi)}(-i)\sqrt{\frac{\omega_{k}}{2}}\left(\hat{a}_{k}e^{ikx}-\hat{a}^{\dagger}_{k}e^{-ikx}\right)\leavevmode\nobreak\ ,$ (77)
where $\omega_{k}=\sqrt{\omega^{2}+k^{2}}$, and where we have made the usual rescaling of the ladder operators by the square-root of the volume, namely $\sqrt{a}\,\hat{a}_{\alpha}\longrightarrow\hat{a}_{k}/{\sqrt{aN}}$, in order to be consistent with $[\hat{a}_{k},\hat{a}_{k^{\prime}}^{\dagger}]=2\pi a\delta(k-k^{\prime})$, versus $[\hat{a}_{\alpha},\hat{a}_{\beta}^{\dagger}]=\delta_{\alpha\beta}$.
A.2 Interacting fields
It is straightforward now to implement an interacting theory consider the Hamiltonian density of scalar QFT in the continuum in its most general form with an arbitrary interaction, $\mathscr{V}$, such that
$\mathscr{H}(x,t)=\frac{1}{2}\left(\pi(x,t)^{2}+\nabla\varphi(x,t)^{2}\right)+\mathscr{V}(\varphi(x,t))\leavevmode\nobreak\ .$ (78)
Following the same discretisation procedure outlined above for the free field theory, on a lattice of $N$ sites the total Hamiltonian for this system becomes,
$H\leavevmode\nobreak\ =\leavevmode\nobreak\ a\sum_{n}\left[\frac{1}{2}\left(\hat{p}_{n}^{2}+\left(\frac{\hat{q}_{\,\overline{n+1}}-\hat{q}_{n}}{a}\right)^{2}\right)+\mathscr{V}_{I}\left(\hat{q}_{n}\right)\right]\leavevmode\nobreak\ .$ (79)
Expanding the terms within the Hamiltonian, we find three kinds of contribution in the total Hamiltonian, the on-site kinetic term, the on-site effective potential and the hopping terms:
$Ha^{-1}=\sum_{n}\left[\frac{1}{2}\hat{p}_{n}^{2}+\mathscr{V}_{\rm eff}(\hat{q}_{n})\right]-\frac{1}{a^{2}}\sum_{n}\hat{q}_{\,\overline{n+1}}\hat{q}_{n}\leavevmode\nobreak\ ,$ (80)
where the on-site effective potential is $\mathscr{V}_{\rm eff}$
$V_{I}(\hat{q}_{n})\leavevmode\nobreak\ =\leavevmode\nobreak\ \frac{1}{a^{2}}\hat{q}_{n}^{2}+\mathscr{V}(\hat{q}_{n})\leavevmode\nobreak\ .$ (81)
In the main body of the paper we set $a=1$, but the overall factor of $a^{-1}$ can be absorbed into our definition of $\delta t$, which can be thought of as expressing the time in lattice units.
Appendix B Sampling displacements and momenta for the initial state of the quantum lattice
In this Appendix we derive the fluctuations required to simulate the false vacuum of the QFT by sampling a classical distribution. To see how this works, first suppose that there were no quantum fluctuations at all in the system, and that we wished to generate all of the fluctuations by sampling. In that case one can proceed as follows. The fluctuations that we wish to reproduce are encompassed by the position and momentum covariance matrices of the QFT vacuum, $C_{q}$ and $C_{p}$, which are $N\times N$ matrices characterising the correlations between sites:
$C_{q}$ $=\langle\hat{q}\hat{q}^{T}\rangle-\langle\hat{q}\rangle\langle\hat{q}^{T}\rangle,$ (82)
$C_{p}$ $=\langle\hat{p}\hat{p}^{T}\rangle-\langle\hat{p}\rangle\langle\hat{p}^{T}\rangle,$ (83)
where $\hat{q}=(\hat{q}_{1},\hat{q}_{2},\ldots,\hat{q}_{N})^{T}$ and similarly for $\hat{p}$. If we Cholesky decompose the covariance matrices as follows,
$C_{q}$ $\leavevmode\nobreak\ =\leavevmode\nobreak\ L_{q}L_{q}^{T},$
$C_{p}$ $\leavevmode\nobreak\ =\leavevmode\nobreak\ L_{p}L_{p}^{T},$ (84)
where $L_{q}$ and $L_{p}$ are lower-triangular matrices, then it is straightforward to show that an ensemble of displacements and momenta of the following form will reproduce the same covariance matrices (where $s$ simply labels the members of the ensemble):
${q}^{(s)}\leavevmode\nobreak$ $=\leavevmode\nobreak\ {q}_{\text{mean}}+L_{q}{\eta}^{(s)}_{q}\leavevmode\nobreak\ ,$
${p}^{(s)}\leavevmode\nobreak$ $=\leavevmode\nobreak\ {p}_{\text{mean}}+L_{p}{\eta}^{(s)}_{p}\leavevmode\nobreak\ .$ (85)
Here $q_{\text{mean}}$ and $p_{\text{mean}}$ are the mean initial displacement and momentum vectors (e.g., $q_{\text{mean}}=\varphi_{\rm fv}$ if starting from the ground state in the false vacuum), and $\eta^{(s)}_{q,p}$ are standard normal random vectors:
${\eta}_{q},{\eta}_{p}\sim\mathcal{N}(\bm{0},\mathbf{I}_{N})\leavevmode\nobreak\ ,$ (86)
where ${\eta}^{(s)}_{q},{\eta}^{(s)}_{p}\in\mathbb{R}^{N}$.
The lattice can then be initialised with position and momentum displacements drawn from this ensemble:
  • Positions: $\{q_{j}^{(s)}\}_{j=1}^{N}$
  • Momenta: $\{p_{j}^{(s)}\}_{j=1}^{N}$
and the dynamical simulation can then proceed as normal.
In the current context, as we are interested mainly in reproducing the missing long-range correlations, it is sufficient to use the free-field SHO approximation around the false vacuum. This is given by [1]
$\langle{\bf q}|\Psi\rangle\leavevmode\nobreak\ \sim\leavevmode\nobreak\ \prod_{\alpha}\left({\omega_{\alpha}}^{-\frac{1}{4}}\leavevmode\nobreak\ e^{-\omega_{\alpha}|q_{\alpha}|^{2}}\right)\leavevmode\nobreak\ ,$ (87)
where the frequencies are given by Eq. (71). Then the computation is considerably simplified. It is straightforward to show that the covariance matrices are
$C_{q,nm}$ $\leavevmode\nobreak\ =\leavevmode\nobreak\ \sum_{\alpha}\frac{1}{2\omega_{\alpha}}U^{\dagger}_{n\alpha}U_{\alpha m}$ (88)
$C_{p,nm}$ $\leavevmode\nobreak\ =\leavevmode\nobreak\ \sum_{\alpha}\frac{\omega_{\alpha}}{2}U^{\dagger}_{n\alpha}U_{\alpha m}\leavevmode\nobreak\ ,$ (89)
where $U_{\alpha m}$ is as given in Eq. (68), and one can then identify
$L_{q,n\alpha}\leavevmode\nobreak\ =\leavevmode\nobreak\ \frac{1}{\sqrt{2\omega_{\alpha}}}U^{\dagger}_{n\alpha}\leavevmode\nobreak\ ;\leavevmode\nobreak\ \leavevmode\nobreak\ L_{p,n\alpha}\leavevmode\nobreak\ =\leavevmode\nobreak\ \sqrt{\frac{\omega_{\alpha}}{2}}\,U^{\dagger}_{n\alpha}\leavevmode\nobreak\ .$ (90)
There is only one subtlety, which is the fact that $q_{n}$ and $p_{n}$ are real. Focussing on the position displacements for the moment, one now has
${q}^{(s)}\leavevmode\nobreak\ =\leavevmode\nobreak\ \varphi_{\rm fv}+\sum_{\alpha=0}^{N-1}\frac{1}{\sqrt{2N\omega_{\alpha}}}e^{-2\pi in\alpha}{\eta}^{(s)}_{q,\alpha}\leavevmode\nobreak\ ,$ (91)
which implies that we must set $\eta^{(s)}_{q,\alpha}=\eta^{(s)}_{q,N-\alpha}$, or equivalently, restoring the lattice index $n$, and selecting from two sets of standard normal vectors $\eta$ and $\eta^{\prime}$, we have
${q}_{n}\leavevmode\nobreak\ =\leavevmode\nobreak\ \varphi_{\rm fv}+\sum_{\alpha=0}^{N-1}\frac{1}{\sqrt{2N\omega_{\alpha}}}\,\cos(2\pi n\alpha/N)\leavevmode\nobreak\ {\eta}_{n,\alpha}\leavevmode\nobreak\ ,$ (92)
and likewise implement initial momenta of the form
$p_{n}\leavevmode\nobreak\ =\leavevmode\nobreak\ \sum_{\alpha=0}^{N-1}\sqrt{\frac{\omega_{\alpha}}{2N}}\,\cos(2\pi n\alpha/N)\leavevmode\nobreak\ {\eta}^{\prime}_{n,\alpha}\leavevmode\nobreak\ .$ (93)
(This constraint is equivalent to requiring that the fluctuations conserve momentum in the system as a whole.) This is the final expression for generating all the initial fluctuations entirely by random sampling of a classical ensemble.
This is not quite what we want for the case at hand, because as we have already discussed the low-momentum fluctuations are all present and correct even when we start the system in an MPS which is just the product of qumode “false vacuum ground states”. What we need to add to this initial system are just those fluctuations of the QFT vacuum which are missing.
[…]
Thus we see that the required covariance matrices encompass the difference between the full contribution and the degenerate contribution from the product of qumode groundstates, namely
$C_{q,nm}$ $\leavevmode\nobreak\ =\leavevmode\nobreak\ \sum_{\alpha}\frac{1}{2}\left(\frac{1}{\omega_{\alpha}}-\frac{1}{\omega}\right)\leavevmode\nobreak\ U^{\dagger}_{n\alpha}U_{\alpha m}$ (94)
$C_{p,nm}$ $\leavevmode\nobreak\ =\leavevmode\nobreak\ \sum_{\alpha}\frac{1}{2}\left(\omega_{\alpha}-\omega\right)\leavevmode\nobreak\ U^{\dagger}_{n\alpha}U_{\alpha m}\leavevmode\nobreak\ .$ (95)
These covariance matrices can be generated by fluctuations around the product of qumode “false vacuum ground states” of the following form:
$\Delta{q}_{n}\leavevmode\nobreak$ $=\leavevmode\nobreak\ \sum_{\alpha=0}^{N-1}\frac{1}{\sqrt{2N}}\sqrt{\left(\frac{1}{\omega}-\frac{1}{\omega_{\alpha}}\right)}\leavevmode\nobreak\ \cos(2\pi n\alpha/N)\leavevmode\nobreak\ {\eta}_{n,\alpha}\leavevmode\nobreak\ ,$
$\Delta p_{n}\leavevmode\nobreak$ $=\leavevmode\nobreak\ \sum_{\alpha=0}^{N-1}\sqrt{\frac{\omega_{\alpha}-\omega}{2N}}\,\cos(2\pi n\alpha/N)\leavevmode\nobreak\ {\eta}^{\prime}_{n,\alpha}\leavevmode\nobreak\ .$ (96)
Thus as anticipated such fluctuations die away for the low-momentum non-relativistic modes in the spectrum, which as per Eq. (71) satisfy $\alpha/aN\ll\omega$. In the infinite lattice limit, these modes do not contribute to qumode fluctuations, which as mentioned is in accord with the intuition that the QFT vacuum coincides with the product of qumode ground states for the low momentum modes. By contrast, the highly relativistic modes have $\omega_{\alpha}\approx k_{\alpha}=2\pi\alpha/aN$, and the weighted sum retains correlations. It is straightforward to show that the contributions from these relativistic modes to fluctuations in the $q_{n}$ grow slowly with the size of the lattice, as $\Delta q_{n}\sim\sqrt{\log N}$.